OpenStudy (anonymous):

"remove parentheses and simplify" 3m^{4} + m-5(m ^{4} + 7)

OpenStudy (tkhunny):

Hint: -5(r+3) = -5r - 15 Let's see what you get.

OpenStudy (anonymous):

im lost

OpenStudy (tkhunny):

That's not good. You need to know the Distributive Property of Multiplication over Addition. Is is usually stated like this: a(b+c) = ab + ac or maybe (b+c)a = ba + ca You MUST know this property and be able to use it. You will need it in your solution: "-5(m ^{4} + 7)"

OpenStudy (anonymous):

thanks